Orange Corp. has two divisions: Fruit and Flower. The following information for the past year is available for each division:Reg 1A Reg 1B Reg 2 Reg 3 Compute each divisions return on investment (ROI) and residual income for last year. (Enter your RReg 1A Reg 1B Reg 2 Reg 3 Determine which manager seems to be performing better. OFruit Division Flower DivisionSuppose Orange is investing in new technology that will increase each divisions operating income by $120,000. The total inveDetermine whether both managers will support the investment. Fruit Division Manager Flower Division Manager
Answer #1

в 1 1-a 2 Particulars Fruit Flower 3 ROI 8.00% 18.00% 4 Residual Income 61875 178750 5 6 1-b Flower department seems to be woAB 1 1-a 2 Particulars Fruit 3 ROI |=165000/2062500 4 Residual Income =165000-(2062500*5%) Flower =247500/1375000 =247500-(13